OpenStudy (anonymous):

Fine each exact value if (π/2)

OpenStudy (anonymous):

EDIT** Find not fine

OpenStudy (anonymous):

note that \(\sin(x-y) =\sin x \cos y - \sin y \cos x\) and \(\cos x =\sqrt{1-\sin^2 x}\)

OpenStudy (anonymous):

do you need further help?

OpenStudy (anonymous):

I really could use this done out so I can compare it with another problem.

OpenStudy (anonymous):

ah. okaay. so that means you've not done this? since sin y is given, and 0<y<(π/2), then \(\cos y =\sqrt{1-(\frac{4}{5})^2}=\frac{3}{5}\) csc x =7, so sin x=1/7 and cos x (from previous reasoning is \(\frac{2\sqrt3}{7}\)) sub all to get sin(x-y)

OpenStudy (anonymous):

Alright thanks. Just needed to compare with another method to see how it played out.

OpenStudy (anonymous):

ah. :) good luck with the other one.

OpenStudy (anonymous):

lol actually there's another way to this. find the angles x and y and subtract.